Circular No. 7621

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001bg IN NGC 2608 G. M. Hurst, Basingstoke, England, reports the discovery by T. Boles,

Coddenham, of an apparent supernova (mag ∼ 14) on an unfiltered CCD frame (limiting mag ∼ 18) taken on May 8.943 UT with a 0.36-m Schmidt-Cassegrain telescope in the course of the U.K. Nova/Supernova Patrol. The object was near the same brightness on a second image obtained by Boles on May 9.9 in rather poor conditions. The new object is located at α = 8^{h}35^{m}18^{s}.86, δ = +28^{o}28′05′′.8 (equinox 2000.0), which is 22′′ east and 19′′ south of the center of NGC 2608. Nothing appears at this location on a master image taken on 2000 Jan. 13.956 (limiting mag ∼ 19) or on Palomar Sky Survey images taken on 1989 Nov. 8 (limiting red mag ∼ 20) and on 1990 Mar. 23 (limiting blue mag ∼ 20). SN 2001bg also appears near mag 14 on an unfiltered CCD image taken on May 9.884 with a 0.15-m reflector by H. Kerner, Fassberg, Germany, at the request of Hurst.

CI AQUILAE B. E. Schaefer, University of Texas at Austin, writes: “The recurrent

nova CI Aql displayed an eclipsing light curve with a range in V of 14.7 to 15.3 from May 2.32 to 3.48 UT in data obtained with the 2.1-m telescope at McDonald Observatory. The normal maximum during quiescence is V = 16.2, so CI Aql is still in its plateau phase after its May 2000 outburst, during which it is predicted to be a bright supersoft x-ray source (Hachisu

and Kato, http://xxx.lanl.gov/abs/astro-ph/0104461), in need of ob-

servations.”

SUPERNOVA 2001bb IN IC 4319 Corrigendum. On IAUC 7620, while Chassagne’s prediscovery obser-

vation was made of SN 2001bb, the actual discovery (cf. IAUC 7614) was by the LOTOSS, not Chassagne.

COMET C/1999 T1 (McNAUGHT-HARTLEY) Visual m_{1} estimates: Jan. 1.28 UT, 7.9 (B. H. Granslo, Malm, Nor-

way, 0.07-m refractor); Feb. 18.08, 7.9 (J. Carvajal, Avila, Spain, 10×70 binoculars); Mar. 12.88, 9.4 (R. J. Bouma, Groningen, The Netherlands, 0.31-m reflector); Apr. 1.12, 11.3 (N. Biver, Poigny, France, 0.26-m reflector); 28.94, 11.4 (Bouma, 0.25-m reflector).

2001 May 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green